OpenStudy (anonymous):

8(x-2)-5(x+4)=20 / x

OpenStudy (anonymous):

sorry that is supposed to be =20+x

sam (.sam.):

\[\begin{array}{l} 8 (x-2)-5 (x+4)=x+20 \\ \text{Expand out terms of the left hand side:} \\ 3 x-36=x+20 \\ \text{Add }(36-x)\text{ to both sides:} \\ 2 x=56 \\\end{array}\] x=?